OpenStudy (anonymous):

Can someone please help me? Given: f(x) = x² and g(x) = 2x + 1, find f[g(x)].

OpenStudy (anonymous):

2x^2 + 1 4x^2 + 1 4x^2 + 4x + 1

OpenStudy (anonymous):

easy find g(x) first, then plug in that answer into the f(x) equation-that probably makes no sense though. I need an x value or do you just want the combined equation?

OpenStudy (anonymous):

If it is just the combined equation it would be this.. \[x ^{2}(2x+1)\]

OpenStudy (anonymous):

Distributive property...

OpenStudy (anonymous):

f[g(x)] = f*x*g = g(x²) = x²(2x + 1). Correct me if I'm wrong please.

OpenStudy (anonymous):

It's multiple choice and the choices are 2x^2 + 1 4x^2 + 1 4x^2 + 4x + 1

OpenStudy (anonymous):

4x^2 + 1. I'm pretty sure. Correct me if I'm wrong.

OpenStudy (anonymous):

Well if x=2 I would plug it into g(x) an get 2(2)+1=5 Now plug 5 into f(x) 5^2=25

OpenStudy (anonymous):

@Cregannn I think you are right

OpenStudy (anonymous):

DID WE HELP YOU @SubaruWRX

OpenStudy (anonymous):

YES THANKS EVERYONE

OpenStudy (anonymous):

NO PROBLEM BUDDY! @SubaruWRX